Please help me answer the following question: x and y represent the measures of complimentary angles y=2(x-15) please solve for x and y
Answer by ikleyn(52779)

.
Complementary angles sum up to 90 degrees.

Hence, x + y = 90.

Then y = 90 -x.

Therefore (!), 90-x = 2*(x-15).

Simplify and solve this equation for "x".

Then find "y".


On complementary, supplementary, vertical angles see the lessons
    - Angles basics
    - Vertical angles
in this site.

Also, you have this free of charge online textbook on Geometry
    - GEOMETRY - YOUR ONLINE TEXTBOOK
in this site.

The referred lessons are the part of this textbook under the topic
"Basics of angles. Supplementary, complementary angles. Vertical angles. Parallel lines".


--------------------
Also, there is this bunch of lessons closely related to this:
    - Solved problems on supplementary and complementary angles
    - Solved problems on angles of a triangle
    - Solved problems on angles of a parallelogram
    - Solved problems on angles of a quadrilateral
    - Solved problems on angles of a polygon
    - Solved problems on missed angle of a polygon
    - Solved problems on angles of a regular polygon

The referred lessons are the part of the same textbook under the topic
"Finding angles of triangles, parallelograms and rectangles".